# Evaluating Language Models on Entity Disambiguation in Tables

### v1

@@ Abstract.
Tables are crucial containers of information, but understanding their meaning may be challenging. Indeed, recently, there has been a focus on Semantic Table Interpretation (STI), i.e.,, the task that involves the semantic annotation of tabular data to disambiguate their meaning. Over the years, there has been a surge in interest in data-driven approaches based on deep learning that have increasingly been combined with heuristic-based approaches. In the last period, the advent of Large Language Models (LLMs) has led to a new category of approaches for table annotation. The interest in this research field, characterised by multiple challenges, has led to a proliferation of approaches employing different techniques. However, these approaches have not been consistently evaluated on a common ground, making evaluation and comparison difficult. This work proposes an extensive evaluation of four state-of-the-art (SOTA) approaches — Alligator (formerly s-elBat), Dagobah, TURL, and TableLlama; the first two belong to the family of heuristic-based algorithms, while the others are respectively encoder-only and decoder-only LLMs. The primary objective is to measure the ability of these approaches to solve the entity disambiguation task, with the ultimate aim of charting new research paths in the field.
@@ 1. Introduction
In particular, we focus on Entity Linking (EL) in tables (CEA, in the STI terminology), which is relevant not only to support table understanding but also to support data transformation, integration and enrichment processes, which are particularly interesting from a data management point of view. The Cell-Entity Annotation (CEA) tasks can be broken down into two sub-tasks: candidate Entity Retrieval (ER), where a set of candidates for each mention is collected and, often, associated with an initial score and rank, and Entity Disambiguation (ED), where the best candidate is selected (and, in some case, a decision whether to link or not is also considered (Avogadro2023, )).
When considering approaches to STI and, especially, CEA, it should be considered that the content and the structure of tables may differ significantly, also depending on application-specific features: column headers may have interpretable labels or be omitted; the number of rows can vary from a dozen (e.g., as typical in tables published on the web or in scientific papers) to hundreds thousand or even millions (e.g., in business data); the cells may include reference to well-known entities (e.g., geographical entities) as well as to specific ones (e.g., biological taxa); tables may come with a rich textual context (e.g., caption or other descriptions in web or scientific documents), or no context at all (e.g., in business data) (Liu2023, ).
A first generation of approaches to CEA has exploited different matching heuristics, traditional machine learning approaches based on engineered features (in the following “feature-based ML”), or a combination of both (Liu2023, ). The International Semantic Web Challenge on Tabular Data to Knowledge Graph Matching (SemTab), at its sixth edition in 2024, is a community-driven effort to compare different approaches systematically with a common experimental setting. The SemTab challenge has pushed different researchers to increase the performance of their approaches and publish datasets for evaluating STI approaches. For example, some effort has been dedicated to optimising entity retrieval, considering the limitations of approaches based on SPARQL queries or Wikidata lookup services (Avogadro2022, ). A few methods from this research community have included embeddings based on LLMs and graphs to support some tasks (Huynh2022, ; Dasoulas2023, ), including CEA, yet at a limited scale.
@@
The increased recognition of the power of LLMs has led to a new generation of Generalistic Table Understanding and Manipulation (GTUM) approaches that support STI (CEA, CTA and CPA) and other tasks (e.g., question answering, schema augmentation, row population and fact verification among the others). The first remarkable example is TURL (Deng2022, ), which is based on an adaptation of BERT (Devlin2019, ) to consider the tabular structure, additionally fine-tuned to execute specific tasks (including ED). The latest of these models, TableLlama (Zhang2023, ), is based on the autoregressive Large Language Model (LLM) Llama 2-7B (Touvron2023-2, ), which is fine-tuned with instruction tuning to perform specific tasks. The approach exploits the power of instruction tuning with large contexts to convert ED (and other tasks) into a prompt-based text generation task. These approaches have been trained and tested on datasets that fit their generalistic ambition and reported as SOTA approaches to the considered tasks.
While STI-related approaches and GTUM approaches support ED, an evaluation of these families of approaches on a common experimental ground is missing. In addition, the evaluation of GTUM approaches has focused on their performance on different tasks rather than drilling down on CEA; as a result, it is unclear to what extent these approaches reach SOTA performance under different settings (e.g., also considering different types of tables). Finally, several aspects of CEA that are important from a data management point of view, e.g., scalability, have not been considered at all. We posit that several of these questions are important to understand the current status of ED algorithms for tables, as well as the challenges to be addressed in the future.
With our study, we want to provide a more detailed analysis of the advantages and disadvantages of GTUM ED approaches based on LLMs, especially when compared to those specifically focusing on CEA developed in the context of STI.
@@
Our analysis mainly considers the SOTA generalistic generative model TableLlama (Zhang2023, ), its predecessor TURL (Deng2022, ), a BERT-based encoder-only transformer, and Alligator (Avogadro2023, ), a recent feature-based Machine Learning (ML) approach. These models represent three different inference mechanisms for ED, sketched in Figure 1, each one associated with a set of expected advantages: TableLlama is expected to exploit implicit knowledge of a large LLM; TURL is expected to be a more efficient generalistic model based on a small LLM fine-tuned specifically for the ED task; Alligator exploits a set of features engineered based on the experience with the SemTab challenge, which is further processed by two neural networks to return confidence scores. Each approach is tested in in-domain, out-of-domain, and moderately out-of-domain settings, as more precisely defined in Section 4.2. In addition, we test the moderately out-of-domain fine-tuning of LLMs-based approaches to test generalisation and adaptation capabilities. Finally, we provide some results with an approach that achieved top performance in previous SemTab challenges [^1] With some limitations due to its excessive execution time and the difficulty of properly replicating the published results despite our best effort..
@@
We remark that this paper aims to provide insights on the behavior and impact of GTUM models on ED rather than introducing a new approach. This is inspired by the large body of similar analyses addressing specific problems in NLP (Wang2023, ; kirk2024understanding, ; zhong2024can, ; zhang2024llmeval, ; kandpal2023llmstruggle, ). In our work, we also develop an evaluation protocol that can be used in future work.
@@
    Test the performance of different genres of STI models when used in combination with a realistic candidate retrieval step, performed by the engineered tool LamAPI (Avogadro2022, );
@@ 2. Related Work
Heuristic, ML and Probabilistic-based Approaches. We refer to (Liu2023, ) for a thorough review of STI approaches and ED approaches proposed therein up to 2022. The ED task in the STI can be performed by applying multiple techniques while focusing on different inherent information: i) similarity, ii) contextual information, iii) ML techniques, and iv) probabilistic models. Often, the disambiguation step involves selecting the winning candidate based on heuristics, like the string similarity between the entity label and mention (Limaye2010, ; Zhang2017, ; Oliveira2019, ; Kruit2019, ; Steenwinckel2019, ; Thawani2019, ; Cremaschi2019, ; Cremaschi2020-1, ; Cremaschi2020-2, ; Huynh2020, ; Tyagi2020, ; Shigapov2020, ; Avogadro2021, ; Abdelmageed2021, ; Huynh2021, ; Chen2022, ; Cremaschi2022, ).
@@
Starting from encoder-based approaches, Ditto (Li2020, ) utilizes Transformer-based language models to perform a slightly different task; in fact, the goal is entity-matching between different tables. Doduo (Suhara2022, ) performs Column-Type Annotation (CTA) and CPA using a pre-trained language model, precisely fine-tuning a BERT model on serialised tabular data. Column (columns-properties) types are predicted using a dense layer followed by an output layer with a size corresponding to the number of column (columns-properties) types. Dagobah SL 2022 (Huynh2022, ) employs an ELECTRA-based (electra2020, ) cross-encoder, a variant of the BERT model, which takes a concatenated input, including table headers and the entity description, and outputs a probability value representing the entity’s likelihood concerning the headers. TorchicTab (Dasoulas2023, ) is composed of two sub-systems: TorchicTab-Heuristic and TorchicTab-Classification, with the latter that utilises Doduo (Suhara2022, ) internally.
@@
[…]
Table 1. Statistics of the datasets used to pre-train the considered approaches
Approach Dataset Entities Source Entity Domain
Alligator
SemTab2021 - R2
SemTab2021 - R3
SemTab2022 - R2 (2T)
SemTab2020 - R4
SemTab2019 - R3
SemTab2019 - R1 (T2D)
Total
47.4K
58.9K
994.9K
667.2K
390.4K
8K
2.16M
Graph Queries
Graph Queries
Graph Queries / Web Tables
Graph Queries
Graph Queries
Graph Queries / Web Tables
Cross
@@ 3. Considered Approaches
We select four approaches representative of different categories of algorithms that solve semantic tasks on tables, and, especially CEA: TableLlama is the first autoregressive LLM that is specifically instruction-tuned on tabular data and reports SOTA results (Zhang2023, ); TURL, cited as previous SOTA in (Zhang2023, ), is a BERT-based encoder-only model performing the three STI tasks, including CEA (Deng2022, ); Alligator is a recent feature-based ML algorithm focusing on CEA, is publicly available, and has been evaluated in settings similar to the moderately-out-of-domain settings discussed in this paper (Avogadro2023, ); Dagobah, a heuristic-based algorithm, has been the winner of various rounds of the SemTab challenge in 2020 (Huynh2020, ), 2021 (Huynh2021, ) and 2022 (Huynh2022, ) and is publicly available.
@@
We remark that all the selected approaches perform all three STI tasks (which is rare and not obvious, especially for the LLM-based approaches), i.e., CTA, CPA and CEA, and we focus only on the CEA one.
@@
TURL (Deng2022, ) (Figure 1) introduces the standard BERT pre-training finetuning paradigm for relational Web tables and is composed of three main modules: i) an embedding layer to convert different components of a table into embeddings, ii) a pre-trained TinyBERT Transformer (Jiao2020, ) with structure-awareness to capture both the textual information and relational knowledge, and iii) a final projection layer for pre-training/fine-tuning objectives. The embedding layer is responsible for embedding the entire table, distinguishing between word embeddings $\mathbf{x}_{w}$, derived from table metadata and cell text (mentions), and entity embeddings $\mathbf{x}_{e}$, representing the unique entities to be linked. The sequence of tokens in $\mathbf{x}_{w}$ and the entity representation $\mathbf{x}_{e}$ are sent to a structure-aware TinyBERT Transformer to obtain a contextualised representation. To inject the table structural information into the contextualised representations, a so-called visibility-matrix $M$ is created, such that entities and text content in the same row or column are visible to each other, except table metadata that are visible to all table components. During pre-training, both the standard Masked Language Model (MLM) and the newly introduced Masked Entity Retrieval (MER) objectives are employed, with the projection layer that is learned to retrieve both the masked tokens and entities [^2] The entities are retrieved from a small candidates set, considering that entity vocabulary could be quite large. During fine-tuning, the model is specialised to address the specific downstream task. Specifically, for CEA, when provided with the sub-table containing all mentions to be linked to an external KG, TURL is fine-tuned to produce a probability distribution over the combined set of candidates for each mention to be linked. This means that TURL lacks the context provided by all the not-to-be-linked cells, without the possibility of abstaining from answering or responding with a NIL. Also, in the original paper, we observe that the best candidate is chosen by a function that compares the best score computed by TURL and the score assigned by the model to the first candidate retrieved by the Wikidata-Lookup service by down-weighting the best model prediction. As in previous comparisons (Zhang2023, ), to evaluate TURL’s performance on the ED task, we consider the model predictions only. Finally, as TURL accepts the whole table as input, computational performance varies depending on the size of the input table.
@@
TableLlama (Zhang2023, ) (Figure 1) employs Llama2 (Touvron2023-2, ) and instruction tuning (wei2022instruction, ) to solve multiple tasks related to tables, e.g., CTA, CPA, CEA and Q&A to name a few. To this end a multi-task instruction tuning dataset named TableInstruct is made available, containing more than 1.24M training tables and 2.6M instances gathered from 14 different table datasets of 11 distinctive tasks. To account for longer tables, exceeding the maximum context length of Llama2 (which is fixed at 4096 tokens), LongLora (Chen2024, ) is used to enlarge the maximum context length to 8192 tokens. In particular, the LLM is prompted with i) an instruction that describes the high-level task to be solved, ii) an input prepended by the [TLE] special character followed by the table metadata, if available, and the serialised table and iii) a question based on the task to be solved by the LLM. In particular, the table starts with the [TAB] special character and is followed first by the table header and then by every row in the table, separated by the special character separator [SEP]. For the CEA task, the question asks the LLM to link a particular mention found in the table against a small set of candidates (maximum 50 candidates) by extracting the correct candidate from the proposed list without, as in TURL, the possibility of not answer or to answer with a NIL. Since the context is fixed to 8192 tokens, compromises must be made to reduce either the length of the set of candidates or the input table.
@@
Alligator (Figure 1) is a feature-based ML approach based on multiple steps: i) data analysis and pre-processing, ii) ER, iii) local feature extraction, iv) local scoring, v) feature enrichment, vi) context-based scoring. The data analysis and pre-processing step converts all cells to lowercase and removes extra spaces and special characters, e.g., underscores (_), to improve the results of the entity retrieval phase. In addition, columns are classified as either L-column (containing literals) or NE-column (containing named-entity mentions). The ER step extracts relevant candidates from the KG using the LamAPI services. The local feature extraction step builds a vector of engineered features for each candidate entity; the vector represents information about the specific candidate (e.g., popularity, number of tokens) or its comparison with the values in the cell and the row (different similarity scores, matches with other cells on the row, matches with the entity description, and so on). Local scoring computes a matching score for each candidate, using a simple deep Neural Network (NN): the NN takes the local feature vectors as input and is trained to convert the features in a normalised matching score. In practice, this step re-ranks all the candidates, considering mainly text similarity and matches against values on the same row. The last steps compute updated scores for the candidate of a given cell by considering the best candidates for other cells on the same column. In practice, Columns-Property Annotation (CPA) and CTA tasks are executed on a best-effort basis to capture, for each candidate, the degree of agreement between its types and properties and the types of properties of other best candidates in the same column. After this feature enrichment step, a simple deep NN similar to the previous one predicts the updated normalised scores for all the candidates [^3] In principle, the approach also estimates a confidence score for each cell to make a decision whether to link the best candidate or not; however, in this paper, we focus on the ED task and consider the candidate with the best score as the output of Alligator.
@@
Dagobah (Huynh2021, ) performs the CEA, CPA and CTA tasks by implementing a multi-step pipeline: i) table pre-processing, ii) ER, iii) candidate pre-scoring, iv) CPA, CTA, and v) CEA. The table pre-processing step involves generating metadata for the table, including orientation detection, header detection, key column detection, and column primitive typing. These tasks help identify the structure and annotation targets, facilitating subsequent annotation steps. ER retrieves relevant candidate entities from the KG for each (entity) cell in the table using Elasticsearch; it considers the primitive types identified during pre-processing and enriches entity information with aliases to increase coverage. Candidate pre-scoring computes a relevance score for each candidate by combining two features: context and literal similarity. Given a cell to disambiguate and a candidate entity, the context feature compares the cell neighbours to the candidate neighbours using a smart procedure. Overall, this method improves the precision of context scoring by incorporating both direct and indirect connections while avoiding overly complex and noisy paths. Then, CPA and CTA modules identify the most suitable relations and types for column pairs and target columns using a majority voting strategy. Finally, the most relevant entity for each table cell is selected by combining pre-scoring with information from CPA and CTA into a final score. The main challenges with Dagobah are its high time complexity and memory usage, making it impractical for processing large tables. This difficulty stems from the expensive task of matching table rows with relations in a KG. Consequently, we limited our evaluation of DAGOBAH to the datasets HT-R1, HT-R2, and WikidataTablesR1 because these datasets contain tables with limited size, as detailed in Section 4.1.
@@
Table 1 lists the datasets used to train the selected approaches; in particular, for TableLlama we use the published model on HugginFace, while we train TURL and Alligator on the same datasets used in the original paper. We use the term ”pre-train” to refer to this main training phase to distinguish it from subsequent fine-tuning (see Section 4.4). Dagobah is not listed because is not based on training (we use the open-source code).

### v2

# Evaluating LLMs on Entity Disambiguation in Tables

@@ Abstract.
Tables are crucial containers of information, but understanding their meaning may be challenging. Over the years, there has been a surge in interest in data-driven approaches based on deep learning that have increasingly been combined with heuristic-based ones. In the last period, the advent of Large Language Models (LLMs) has led to a new category of approaches for table annotation. However, these approaches have not been consistently evaluated on a common ground, making evaluation and comparison difficult. This work proposes an extensive evaluation of four STI state-of-the-art (SOTA) approaches — Alligator (formerly s-elBat), Dagobah, TURL, and TableLlama; the first two belong to the family of heuristic-based algorithms, while the others are respectively encoder-only and decoder-only LLMs. We also include in the evaluation both GPT-4o and GPT-4o-mini, since they excel in various public benchmarks. The primary objective is to measure the ability of these approaches to solve the entity disambiguation task with respect to both the performance achieved on a common-ground evaluation setting and the computational and cost requirements involved, with the ultimate aim of charting new research paths in the field.
@@ 1. Introduction
In particular, we focus on Entity Linking (EL) in tables (CEA, in the STI terminology), which is relevant not only to support table understanding but also to support data transformation, integration and enrichment processes, which are particularly interesting from a data management point of view. The Cell-Entity Annotation (CEA) task can be broken down into two sub-tasks: candidate Entity Retrieval (ER), where a set of candidates for each mention is collected and, often, associated with an initial score and rank, and Entity Disambiguation (ED), where the best candidate is selected (and, in some case, a decision whether to link or not is also considered (Avogadro2023, )). When considering approaches to STI and, especially, CEA, it should be considered that the content and the structure of tables may differ significantly, also depending on application-specific features: column headers may have interpretable labels or be omitted; the number of rows can vary from a dozen (e.g., as typical in tables published on the web or in scientific papers) to hundreds thousand or even millions (e.g., in business data); the cells may include reference to well-known entities (e.g., geographical entities) as well as to specific ones (e.g., biological taxa); tables may come with a rich textual context (e.g., caption or other descriptions in web or scientific documents), or no context at all (e.g., in business data) (Liu2023, ).
A first generation of approaches for CEA has exploited different matching heuristics, traditional machine learning approaches based on engineered features (in the following “feature-based ML”), or a combination of both (Liu2023, ).
@@
The increased recognition of the power of LLMs has led to a new generation of Generalistic Table Understanding and Manipulation (GTUM) approaches that support STI (CEA, CTA, and CPA) and other tasks (e.g., question answering, schema augmentation, row population, and fact verification among the others). The first remarkable example is TURL (Deng2022, ), which is based on an adaptation of BERT (Devlin2019, ) to consider the tabular structure, additionally fine-tuned to execute specific tasks (including ED). The latest of these models, TableLlama (Zhang2023, ), is based on the autoregressive Large Language Model (LLM) Llama 2-7B (Touvron2023-2, ), which is fine-tuned with instruction tuning(wei2022instruction, ) to perform specific tasks. These approaches have been trained and tested on datasets that fit their generalistic ambition and reported as SOTA approaches for the considered tasks.
While both STI-related and GTUM approaches facilitate entity disambiguation (ED), a comparative evaluation of these two method families on a common experimental foundation has yet to be conducted. STI-related methods have primarily been assessed within the International Semantic Web Challenge on Tabular Data to Knowledge Graph Matching (SemTab), where most tables are synthetically generated by querying knowledge graphs (KGs), closely resembling tables from industrial or corporate environments. These tables, however, lack the rich semantic context of web tables found on platforms like Wikipedia. In contrast, GTUM approaches have been evaluated mainly on web-based tables, which are semantically rich and crafted by humans for human use. Consequently, it remains unclear how these approaches perform relative to the state of the art (SOTA) under different conditions. Finally, several aspects of CEA that are important from a data management point of view, e.g., scalability, have not been considered. We posit that several of these questions are important to understand the current status of ED algorithms for tables, as well as the challenges to be addressed in the future.
@@
With our study, we want to provide a more detailed analysis of the advantages and disadvantages of GTUM ED approaches based on LLMs, especially when compared to those specifically focusing on CEA developed in the context of STI. Our analysis mainly considers the SOTA generalistic generative model TableLlama (Zhang2023, ), its predecessor TURL (Deng2022, ), a BERT-based encoder-only transformer, and Alligator (Avogadro2023, ), a recent feature-based Machine Learning (ML) approach. These models represent three different inference mechanisms for ED, sketched in Figure 1, each one associated with a set of expected advantages: TableLlama is expected to exploit implicit knowledge of a large LLM; TURL is expected to be a more efficient generalistic model based on a small LLM fine-tuned specifically for the ED task; Alligator exploits a set of features engineered based on the experience with the SemTab challenge, which is further processed by two small neural networks to return confidence scores. Each approach is tested in in-domain, out-of-domain, and moderately out-of-domain settings, as more precisely defined in Section 4.2. In addition, we test the moderately out-of-domain fine-tuning to test generalisation and adaptation capabilities of TURL, TableLlama and Alligator. Finally, we provide the results obtained by Dagobah, an approach that achieved top performance in previous SemTab challenges [^1] With some limitations due to its excessive execution time and the difficulty of properly replicating the published results despite our best effort..
@@
We remark that this paper aims to provide insights on the behavior and impact of GTUM models on ED rather than introducing a new approach. This is inspired by the large body of similar analyses addressing specific problems in NLP (Wang2023, ; kirk2024understanding, ; zhong2024can, ; zhang2024llmeval, ; kandpal2023llmstruggle, ). In our work, we also develop an evaluation protocol that can be used in future research.
@@
    Test the performance of different genres of STI models when used in combination with a realistic candidate retrieval method;
@@ 2. Related Work
Heuristic, ML and Probabilistic-based Approaches. We refer to (Liu2023, ) for a thorough review of STI approaches and ED approaches proposed therein up to 2022. The ED task in STI can be performed by applying multiple techniques while focusing on different inherent information: i) similarity, ii) contextual information, iii) ML techniques, and iv) probabilistic models. Often, the disambiguation step involves selecting the winning candidate based on heuristics, like the string similarity between the entity label and mention (Limaye2010, ; Zhang2017, ; Oliveira2019, ; Kruit2019, ; Steenwinckel2019, ; Thawani2019, ; Cremaschi2019, ; Cremaschi2020-1, ; Cremaschi2020-2, ; Huynh2020, ; Tyagi2020, ; Shigapov2020, ; Avogadro2021, ; Abdelmageed2021, ; Huynh2021, ; Chen2022, ; Cremaschi2022, ).
@@
Starting from encoder-based approaches, Ditto (Li2020, ) utilizes Transformer-based language models to perform a slightly different task, i.e., entity-matching between different tables. Doduo (Suhara2022, ) performs Column-Type Annotation (CTA) and CPA using a pre-trained language model, precisely fine-tuning a BERT model on serialised tabular data. Column (columns-properties) types are predicted using a dense layer followed by an output layer with a size corresponding to the number of column (columns-properties) types. Dagobah SL 2022 (Huynh2022, ) employs an ELECTRA-based (electra2020, ) cross-encoder, a variant of the BERT model, which takes a concatenated input, including table headers and the entity description, and outputs a probability value representing the entity’s likelihood concerning the headers. TorchicTab (Dasoulas2023, ) is composed of two sub-systems: TorchicTab-Heuristic and TorchicTab-Classification, with the latter that utilises Doduo (Suhara2022, ) internally.
@@
Table 1. Statistics of the datasets used to pre-train the selected approaches
Approach Dataset Entities Source Entities Domain
Alligator (Avogadro2023, ) SemTab2021 - R2 47.4K Graph Queries Cross
SemTab2021 - R3 58.9K Graph Queries
SemTab2022 - R2 (2T) 994.9K Graph Queries / Web Tables
SemTab2020 - R4 667.2K Graph Queries
SemTab2019 - R3 390.4K Graph Queries
SemTab2019 - R1 (T2D) 8K Graph Queries / Web Tables
Total 2.16M
@@ 3. Selected Approaches
We select four STI approaches representative of different categories of algorithms that solve semantic tasks on tables, and, especially CEA: TableLlama is the first autoregressive LLM that is specifically instruction-tuned on tabular data and reports SOTA results (Zhang2023, ); TURL, cited as previous SOTA in (Zhang2023, ), is a BERT-based encoder-only model performing the three STI tasks, including CEA (Deng2022, ); Alligator is a recent feature-based ML algorithm focusing on CEA, is publicly available, and has been evaluated in settings similar to the moderately-out-of-domain settings discussed in this paper (Avogadro2023, ); Dagobah, a heuristic-based algorithm, has been the winner of various rounds of the SemTab challenge in 2020 (Huynh2020, ), 2021 (Huynh2021, ) and 2022 (Huynh2022, ) and is publicly available.
@@
While we focus our study only on the CEA task, we remark that all the selected approaches perform all three STI tasks (which is rare and not obvious, especially for the LLM-based approaches), i.e., CTA, CPA and CEA.
@@
TURL (Deng2022, ) introduces the standard BERT pre-training/finetuning paradigm for relational Web tables and is composed of three main modules: i) an embedding layer to convert different components of a table into embeddings, ii) a pre-trained TinyBERT Transformer (Jiao2020, ) with structure-awareness to capture both the textual information and relational knowledge, and iii) a final projection layer for pre-training/fine-tuning objectives. The embedding layer is responsible for embedding the entire table, distinguishing between word embeddings $\mathbf{x}_{w}$, derived from table metadata and cell text (mentions), and entity embeddings $\mathbf{x}_{e}$, representing the unique entities to be linked. The sequence of tokens in $\mathbf{x}_{w}$ and the entity representation $\mathbf{x}_{e}$ are sent to a structure-aware TinyBERT Transformer to obtain a contextualised representation. To inject the table structural information into the contextualised representations, a so-called visibility-matrix $M$ is created, such that entities and text content in the same row or column are visible to each other, except table metadata that are visible to all table components. During pre-training, both the standard Masked Language Model (MLM) and the newly introduced Masked Entity Retrieval (MER) objectives are employed, with the projection layer that is learned to retrieve both the masked tokens and entities [^2] The entities are retrieved from a small candidates set, considering that entity vocabulary could be quite large. During fine-tuning, the model is specialised to address the specific downstream task. Specifically, for CEA, when provided with the sub-table containing all mentions to be linked to an external KG, TURL is fine-tuned to produce a probability distribution over the combined set of candidates for each mention to be linked. This means that TURL lacks the context provided by all the not-to-be-linked cells, without the possibility of abstaining from answering or responding with a NIL. Also, in the original paper, we observe that the best candidate is chosen by a function that compares the best score computed by TURL and the score assigned by the model to the first candidate retrieved by the Wikidata-Lookup service by down-weighting the best model prediction. As in previous comparisons (Zhang2023, ), to evaluate TURL’s performance on the ED task, we consider the model predictions only. Finally, as TURL accepts the whole table as input, computational performance varies depending on the size of the input table.
@@
TableLlama (Zhang2023, ) employs Llama2 (Touvron2023-2, ) and instruction tuning (wei2022instruction, ) to solve multiple tasks related to tables, e.g., CTA, CPA, CEA and Q&A to name a few. To this end a multi-task instruction tuning dataset named TableInstruct is made available, containing more than 1.24M training tables and 2.6M instances gathered from 14 different table datasets of 11 distinctive tasks. To account for longer tables, exceeding the maximum context length of Llama2 (which is fixed at 4096 tokens), LongLora (Chen2024, ) is used to enlarge the maximum context length to 8192 tokens. In particular, the LLM is prompted with i) an instruction that describes the high-level task to be solved, ii) an input prepended by the [TLE] special character followed by the table metadata, if available, and the serialised table and iii) a question based on the task to be solved by the LLM. In particular, the table starts with the [TAB] special character and is followed first by the table header and then by every row in the table, separated by the special character separator [SEP]. For the CEA task, the question asks the LLM to link a particular mention found in the table against a small set of candidates (maximum 50 candidates) by extracting the correct candidate from the proposed list without, as in TURL, the possibility of not answer or to answer with a NIL. Since the context is fixed to 8192 tokens, compromises must be made to reduce either the length of the set of candidates or the input table.
Table 2. Comparison of TURL, TableLlama, and Alligator for the CEA task.
TURL (Deng2022, ) TableLlama (Zhang2023, ) Alligator (Avogadro2023, )
Model TinyBERT (Jiao2020, ) equipped with table structure-awareness (visibility matrix) Llama2 with LongLora (Chen2024, ) for extended context Feature-based ML approach with two Neural Network (NN)s for scoring
Input Table comprised of only the mentions to be linked, optional metadata and candidates for every mentions Serialized table with instruction, optional metadata, mention to be linked and the list of candidates, limited to 8192 tokens Whole table (candidates are retrieved on-the-fly thanks to LamAPI)
Objective Pre-training with MLM and MER; task-specific fine-tuning on CEA Multi-task instruction tuning for CEA, CTA, CPA, and other table-related tasks Feature-based binary cross-entropy with negatives sampling: ${p(y_{m}=1|F_{m})\geq p(y_{m}=0|F_{w};w\in\mathcal{N}(m))}$ [^3] Given a mention $m$ and its correct associated entity in the knowledge graph, $p(y_{m}=1\mid F_{m})$ represents the probability that the features $F_{m}$ correctly identify the correct candidate. For each incorrect candidate $w$ in the set $\mathcal{N}(m)$, $p(y_{m}=0\mid F_{w})$ represents the probability that the features $F_{w}$ correctly indicate that $w$ is not the correct candidate.
Output Probability distribution over candidates (Auto-regressively) Extracts candidate directly from list Score in $[0,1]$ for every candidate
Limitations Input table limited to mentions to be linked, no NIL-awareness Limited to 8192 tokens, no score over candidates, no NIL-awareness No semantic-related features, biased towards KG matching
@@
Alligator is a feature-based ML approach based on multiple steps: i) data analysis and pre-processing, ii) ER, iii) local feature extraction, iv) local scoring, v) feature enrichment, vi) context-based scoring. The data analysis and pre-processing step applies standard pre-processing transformations (e.g., lowercase, space removing), with columns classified as either L-column (containing literals) or NE-column (containing named-entity mentions). The ER step extracts relevant candidates from the KG using the LamAPI services. The local feature extraction step builds a vector of engineered features for each candidate entity; the vector represents information about the specific candidate (e.g., its KG popularity, number of tokens and the syntactical similarity between the candidate label and the mention) or its comparison with the row cells content (e.g., trying to match related literals with the row content). Local scoring computes a matching score for each candidate, using a simple deep NN (denoted PN in (Avogadro2023, )): the NN takes the local feature vectors as input and is trained to convert the features into a normalised matching score. In practice, this step re-ranks all the candidates, considering mainly text similarity and matches against values on the same row. The last steps compute updated scores for the candidate of a given mention by considering the best candidates for other cells on the same column. In practice, Columns-Property Annotation (CPA) and CTA tasks are executed on a best-effort basis to capture, for each candidate, the degree of agreement between its types and properties and the types of properties of other best candidates in the same column. After this feature enrichment step, a simple deep NN (denoted RN in (Avogadro2023, )) similar to the previous one predicts the updated normalised scores for all the candidates [^4] In principle, the approach also estimates a confidence score for each cell to make a decision whether to link the best candidate or not; however, in this paper, we focus on the ED task and consider the candidate with the best score as the output of Alligator.
@@
Dagobah (Huynh2021, ) performs the CEA, CPA and CTA tasks by implementing a multi-step pipeline: i) table pre-processing, ii) ER, iii) candidate pre-scoring, iv) CPA, CTA, and v) CEA. The table pre-processing step involves estimating table metadata, e.g., header detection, and column type estimation. These tasks help identify the structure and annotation targets, facilitating subsequent annotation steps. ER retrieves relevant candidate entities from the KG for each mention in the table using Elasticsearch; it considers the primitive types identified during pre-processing and enriches entity information with aliases to increase coverage. Then, the candidate pre-scoring computes a relevance score for each candidate by combining two features: context and literal similarity. Given a mention to disambiguate and a candidate entity, the context feature compares the mention neighbors to the candidate neighbors with respect to the underline KG. Finally, the most relevant entity for each mention is selected by combining pre-scoring with information from CPA and CTA into a final score. The main challenges with Dagobah are its high time complexity and memory usage, making it impractical for processing large tables. This difficulty stems from the expensive task of matching table rows with relations in a KG. Consequently, we limited our evaluation of Dagobah to the datasets HT-R1, HT-R2, and WikidataTables-R1 because these datasets contain tables with limited size, as detailed in Section 4.1.
@@
Table 1 lists the datasets used to train the selected approaches; in particular, for TableLlama we use the published model on HugginFace, while we train TURL and Alligator on the same datasets used in the original paper. We use the term ”pre-train” to refer to this main training phase to distinguish it from subsequent fine-tuning (cfr. Section 4.4). Dagobah is not listed because is not based on training.
Dataset Split Tables Cols (min |||| max |||| $\bar{x}$) Rows (min |||| max |||| $\bar{x}$) Entities
Train† 91 1 |||| 8 |||| 4,86 5 |||| 369 |||| 98,61 14 674
SemTab2022 2T-red Test† 26 1 |||| 8 |||| 4,65 7 |||| 264 |||| 74,58 4 691
Train 3 691 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,68 26 189
SemTab2022 HT-R1 Test 200 2 |||| 5 |||| 2,59 4 |||| 8 |||| 5,74 1 406
Train† 4 344 2 |||| 5 |||| 2,56 4 |||| 8 |||| 5,57 20 407
SemTab2022 HT-R2 Test 426 2 |||| 5 |||| 2,53 4 |||| 8 |||| 5,56 1 829
Train 9 917 2 |||| 4 |||| 2,51 3 |||| 11 |||| 5,65 64 542
SemTab2023 WikidataTables-R1 Test 500 2 |||| 4 |||| 2,46 3 |||| 11 |||| 6,95 4 247
TURL-120k Train 13 061 1 |||| 43 |||| 5,43 1 |||| 624 |||| 12,98 120 000
TURL-2k-red Test 1 295 1 |||| 14 |||| 1,03 6 |||| 257 |||| 32,95 1 801